OpenStudy (anonymous):

If the sides of a square are lengthened by 5 cm, the area becomes 81 cm^2. Find the length of a side of the original square.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\sqrt[20]{(-9)^{20}} \] need to simplify my answer.

OpenStudy (anonymous):

let x be the original length than the new side will be (x+5) Area of a square is side^2 (x+5)^2=81 x^2+10x-56=0 (x+14)(x-4)=0 x=-14 or x=4 clearly -14 cannot be a length so the original length was 4